# Removing Reflections from RAW Photos

### v2

@@ 4 Reflection removal
Our system removes reflections from RAW images, $m$, in linear RGB color (ACR Step 6) with an optional context image $c$ that is white balanced like $m$. See Func. 1. Both $m$ and $c$ share a scene-referred color space, which aids removal; RGB supports pre-trained perceptual losses. We predict $t$ and $r$ in linear RGB, and store inference outputs by inverting ACR steps 36 to produce new RAW images.
Our system uses two models. A base model operates on $(m,c)$ at $256^{2}$ pixels to predict $(t,r)$ (rectangular images are tiled and linearly blended). These outputs are passed to an upsampler that is applied at each level of a Gaussian pyramid. The models apply these conceptual steps: 1) inputs $m$ are projected to a higher, $d$-dimensional space; 2) nonlinear, “semantic” per-pixel features are computed; 3) components $t$ and $r$ are identified in $d$; 4) output images are rendered. We implement these steps differently in each model.
@@ 4.1 Base model
The base model is illustrated in Fig. 2. A multi-scale feature backbone projects $m$ into a linear, high dimensional space and computes semantic features (labeled P-Net). The multi-resolution semantic features are fused (labeled F-Net) with a feature pyramid network (FPN) at the input resolution. We use EfficientNet as the backbone [44] at $256^{2}$ pixels and fuse features with a BiFPN pyramid [45, 53].
The context image $c$, is processed identically to $m$. Its lowest resolution FPN features are used to predict affines that progressively modify the FPN features of $m$ using conv-mod-deconv operations ala StyleGAN [23]. Modulation is per-channel because $c$ does not share identical content with $m$. Conceptually, modulation gives the model additional capacity to identify $r$ in its features. A finishing module further identifies and renders $t,r$ (it is the head in [58]). We predict $(t,r)$ independently, rather than enforcing $t+r=m$, to decouple failures. Training uses the losses of [58] with improvements to the adversary and gradient terms. Crucially, training is end-to-end from random weights. See Sec. E.1.
Figure 3: Upsampler at one pyramid level.
@@ 4.2 Upsampler
The upsampler, Fig. 3, is iteratively applied over a Gaussian pyramid. It first projects the low- and high-resolution images $(m,r,t)$, and $M$ into a high dimensional space $\phi$ with a convolutional backbone. Critically, batch norms are omitted so the same transform is applied to $(m,r,t)$ and $M$.
The upsampler first identifies features $\phi_{t}$, $\phi_{r}$ in $\phi_{m}$ via feature masks. This matching process uses products of features: when activations match, their product can be large regardless of sign, whereas summation yields large activations if either input is large. We generalize this idea in a mask prediction module, Fig. 3 (bottom), which predicts affines to match and scale activations before a sigmoid is applied. Two per-pixel, per-channel masks are predicted, $\mathbb{I}_{t}$, $\mathbb{I}_{r}$. Errors are corrected by a joint predictor (a per-channel, per-pixel MLP; the affines are predicted similarly; see Sec. E.2 for details). Masks, $(\mathbb{I}_{t}$, $\mathbb{I}_{r})$, are resampled 2×2$\times$2 × and multiplied with $\phi_{M}$ to project its features into subspaces for $T$, $R$. By resampling masks, not features, sharp features are preserved. This key step assumes the identity $(\mathbb{I}_{t}$, $\mathbb{I}_{r})$ of the component to which each feature belongs is low frequency. Errors are corrected with finishing convolutions, which render $T,R$.
Training uses a cycle-consistency loss on $(t,r)$, and the losses of [36], but with LPIPS [57] rather than contextual loss [35] because images are aligned. The model is pre-trained from random weights on ground truth, and fine tuned with the base model. See Sec. E.2 for training details.
@@ 5 Results
We compare the simulation, base model, and upsampler to prior methods.
@@ 5.1 Reflection simulation
Figure 4: The importance of synthesizing training data (top row) from linear domain images (middle row), with comparisons to prior work. (a) Photometrically accurate illuminant colors are simulated by mixing before white balancing; mixing 8-bit white balanced images is much different. (b) Mixing in scene-referred linear units produces reflections that are strong in the shadows, but transparent in the highlights. (prior work) These effects are visibly incorrect in prior methods, which blend $8$-bit tone mapped images [51, 11]. (bottom) Real and simulated examples are shuffled together. For each real image, a similar synthetic reflection was manually found in the dataset. Real images were not captured to match known examples; these qualitative matches exist because the dataset size exceeds $10^{6}$ (even numbered images are synthetic) .
@@
Source images are drawn from MIT5K [8], RAISE [10], and Laval Indoors [14] for a total of $12$,$803$ RAW images and $2$,$233$ scene-referred Image-Based Lighting (IBL) panoramas. The $360^{\circ}$ IBLs are equivalent to about $12$,$367$ indoor RAW images because we simulate random cameras (the expected FOV is $65^{\circ}$, Sec. B.2, B.4). Images are grouped into $10$,$547$ outdoor and (in mathematical expectation) $14$,$623$ indoor images to create pairs $(i,j)\in\mathcal{D}$, Sec. D.2. Indoor and outdoor images are split into train, validation, and test sets so no images are shared among the sets. We use $80\%$ for training and $15\%$ for test.
@@
The number of examples $(m,t,r)$ is amplified by randomization in the geometric simulation (Appendix B, D). We search $10^{8}$ examples for useful $m$. After culling, about $10^{7}$ mixtures remain, and we rendered $10\%$ at $256^{2}$ and $2048^{2}$ pixels to train the base and upsampler models. The $256^{2}$ pixel dataset has $1$,$241$,$091$ for training, $46$,$121$ for test, and $8$,$991$ for validation; the $2048^{2}$ dataset has $1$,$079$,$631$, $39$,$916$, and $7$,$448$ (smaller due to resolution constraints).
@@
Fig. 4 shows results of mixing scene-referred images: (a) correct illuminant colors and (b) correct reflection visibility. We linearly blended $8$-bit tone-mapped images for comparison, and compare to prior works (see caption).
@@
Discussion. Prior works mix $8$-bit tone-mapped images, and the results are qualitatively unrealistic. In particular, their simulated reflections overpower the highlights, and are not powerful enough in the shadows, which are boosted by tone mapping. By contrast, in our simulations light from two scenes is mixed linearly and equally without tone-mapping. This accurate mixing allows our system to generalize better to new scenes, and gives us SOTA performance without the need to train on real mixture images. Furthermore, by synthesizing photometrically and geometrically plausible mixture images, we naturally encode priors on the appearance of reflections in real-world scenes. For example, indoor light is typically relatively weak, so reflections from indoor scenes are typically of regions that are close to light sources (ceilings, etc.). These create small reflection regions that often look yellow when they appear over outdoor scenes, due to typical indoor illuminant colors. Conversely, outdoor light is powerful enough to bounce off diffuse objects, and have enough strength to create reflections of whole scenes that can be colorful, or blue in white balances due to the color of outdoor illuminants. Of course, at nighttime, whole indoor scenes can reflect over cityscapes, etc. These kinds of priors are apparent in consumer photos (see Figs. 1, 9, 9, A12, A13, A14). Lastly, like prior works we pair indoor/outdoor photos, which permits unlikely but plausible pairings such as bathrooms and beaches. Future work can remove such unlikely pairings if they prove unhelpful.
@@ 5.2 Base reflection removal
We trained three base models, one with and two without context $c$. To omit $c$, we remove the modulated merges (see Fig. 2), which removes generally useful model capacity. As a second alternative, the base architecture is not changed, and the model is trained and tested with random $c$. We use this latter randomized approach for ablation. A third alternative, setting $c\leftarrow\mathrm{const}$, forces the backbones to handle both natural $m$ and unnatural $c$, and yields unreliable performance.
Figure 5: Reflection removal results. Outputs at $256\times 341$ were upsampled to $2048\times 2731$ using our upsampler. Base outputs are inset.
@@
Control $21.74$ $0.858$ $12.48$ $0.490$
Ours+ctx $\mathbf{33.23}$ $\mathbf{0.957}$ $\mathbf{30.17}$ $\mathbf{0.882}$
Ours $32.15$ $0.952$ $29.18$ $0.867$
DSRNet [20] $28.98$ $0.926$ $23.99$ $0.755$
Zhang [58] $26.23$ $0.899$ $22.78$ $0.615$
CoRRN [47] $22.75$ $0.867$ $18.31$ $0.604$
Abl. Ours+rnd $32.42$ $0.951$ $29.29$ $0.867$
Ours+rac $33.20$ $0.957$ $30.20$ $0.883$
Table 1: Base model performance. Prior methods were trained on our data. (Control) shows the average degradation to $t$, $r$. Ablations: (rac) the GT reflection is used as context; (rnd) random images are used as context. The context model (ctx) is better without random context ($p<1.7e$-$11$).
Method PSNRt SSIMt PSNRr SSIMr
GT Control $19.50$ $0.863$ $13.13$ $0.645$
Ours $\mathbf{47.77}$ $\mathbf{0.988}$ $\mathbf{45.93}$ $\mathbf{0.982}$
Ours-NM $43.29$ $0.981$ $43.99$ $0.969$
VDSR+C $42.24$ $0.979$ $38.32$ $0.938$
VDSR $40.74$ $0.974$ $38.30$ $0.939$
E2E Ours $\mathbf{30.62}$ $\mathbf{0.952}$ $\mathbf{28.53}$ $\mathbf{0.907}$
VDSR+C $30.27$ $0.945$ $27.74$ $0.887$
@@
Table 2: Upsampler performance at $2048^{2}$, upsampling ground truth (GT), and using the base model for end-to-end operation (E2E). Images are upsampled from $256^{2}$ to $2048^{2}$ using our method and V-DESIRR [36] (VDSR) with and without cycle consistency (+C), which improves VDSR for $\mathrm{T}$ ($p<1e$-$12$).
@@
We compare to Zhang et al. [58], DSRNet [20], and CoRRN [47] by training their models on our dataset. Recall that our model uses the same losses and network head as Zhang et al. to simplify comparison to prior work. Tab. 2 shows that all methods improve images relative to the average degradation to $t,r$ (control). Our models outperform prior works (ours+ctx, ours). We ablate by using random $c$ (ours+rnd), which degrades performance, and this is statistically significant. Removing operations that use $c$ (ours) does not degrade performance compared to training with, and using, random $c$ (differences are not significant), which suggests the contextual model does not learn better priors with its additional capacity, but rather leverages the content of $c$. Ablating further, using the reflection as the context ($c=r$) at test time only does not improve the contextual model performance (ours+rac), which suggests the model has learned to be appropriately sensitive to $c$: the model does not require $c$ and $r$ to match, and it is robust to their differences because it is trained with disjoint crops ($c$, $r$).
@@
For visual comparison, we captured [^3] We thank Florian Kainz for his help capturing these photos. ground truth reflections in three common cases: looking out of a house, looking into a display case, and photographing artwork Fig. 9 and Fig. A12. We dereflected with Zhang, DSRNet (retrained, RT), and our models at $256\times 384$ (inset images) and upsampled them to $2048\times 2731$ (see next section). The empirical SSIM values (lowercase t, r) are commensurate with test performance (Tab. 2). In Fig. 9 our contextual model separates the reflection, but without context our model attributes the colors in the umbrella with a reflected object. Prior works perform quantitatively worse.
@@
In Figs. 1, 5, 9, A13, A14, and A15 we compare results on photos in-the-wild from cameras that were not used to construct the synthetic dataset. We also compare the published $8$-bit models of Zhang and DSRNet (excepting Fig. 5). The 8-bit models do not recover most $t$ or $r$, but generalize better when re-trained on our data (Figs. 9, A13, A14). These methods do not however recover $r$ well, which is needed for aesthetic and error corrections (see Sec. 5.4).
Our model has $2.5$M parameters, while DSRNet has $125$M. Inference at $256\times 341$, takes $0.96\mathrm{s}$/$1.04\mathrm{s}$ on a $2021$ M1 MacBook Pro ($32$Gb) and iPhone $14$ Pro.
@@
Discussion. Our models recover $(t,r)$ in diverse real-world cases including museums, nature, shopping, a mid-day city, artwork, etc. (Figs. 1, 5, 9, 9, A12, A13, A14, A15). In Fig. 1 the contextual model yields more correct and uniform color on the Egyptian tablet by reducing ambiguity about the color of the reflection scene, which is typical (compare to inset w/o context). Failures occur when $t$ or $r$ is bright, and pushes the other into the noise floor, effectively saturating it to black. As reflections strengthen, the removal problem transitions from unmixing to hole filling. When a single color channel saturates, RAW images can preserve enough information to recover the lost content. Nonetheless, systems must address both sub-problems because users typically cannot control the strength of reflections.
Figure 6: Upsampler performance. We upsample the ground truth image from $256\times 384$ to $2048\times 3072$ using our method and V-DESIRR [36]. The latter creates artifacts. Our model eliminates these artifacts while using $29\%$ fewer parameters.
@@
Lastly, errors can occur when textured regions of $t$ and $r$ overlap, as in Fig. 1 where a stone wall overlaps with the dress of the subject. Color differences can help separate complex textures, as in Fig. 9 where the reflection of a painting is successfully separated from complex tree textures. Without such differences, overlapping textures corrupt $t$, and models must repair or hallucinate the missing content as they might also do in saturated regions.
Figure 7: Predicting the reflection enables aesthetic editing and error correction. Here we edited the reflection color, and re-inserted some to show what the dog is looking at.

### v3

@@ 4 Reflection removal
Our system removes reflections from RAW images, $m$, in linear RGB color (ACR Step 6) with an optional context image $c$ that is white balanced like $m$ (see Func. 3.3). Both $m$ and $c$ share a scene-referred color space, which aids removal; RGB supports pre-trained perceptual losses. We predict $t$ and $r$ in linear RGB, and store inference outputs by inverting ACR steps 36 to produce new RAW images.
Figure 2: System
Our system uses two models, Fig. 2. A base model uses $m,c$ at $256^{2}$ pixels to predict $t$, $r$ (rectangular images are tiled); $t,r$ are then upsampled using a Gaussian pyramid.
@@ 4.1 Base model
The base model is in Fig. S4 due to space limits. A multi-scale backbone projects $m$ into a high dimensional space and computes semantic features (labeled P-Net). These features are fused (labeled F-Net) with a feature pyramid network (FPN) at the input resolution. The backbone is EfficientNet [43] at $256^{2}$ pixels, and fusion uses a BiFPN pyramid [44, 54].
The context image $c$, is processed identically to $m$. Its low-resolution FPN features are used to predict affines that modify the FPN features of $m$ using conv-mod-deconv operations ala StyleGAN [23]. Modulation is per-channel because $c$ does not share identical content with $m$. Conceptually, modulation gives the model additional capacity to identify $r$ within the features of $m$. A finishing module further identifies and renders $t,r$ (it’s the head in [59]). We predict $(t,r)$ independently, rather than enforcing $t+r=m$, to decouple failures. Training uses the losses of [59] with improvements to the adversary and gradient terms. Crucially, training is end-to-end from random weights. See Sec. E.1.
@@ 4.2 Upsampler
The upsampler is shown in Fig. S5 due to space limits. Upsampling is performed iteratively over a Gaussian pyramid (see Fig. 2), as summarized below. Details are in Sec. E.2.
Briefly, the upsampler first projects the low- and high-resolution images $(m,r,t)$, and $M$ into a high dimensional space $\phi$ using a convolutional backbone. The upsampler then matches low resolution features $\phi_{t}$, $\phi_{r}$ to $\phi_{m}$ to create masks that identify the features of $t$, $r$ within $m$. This matching process uses products of features: when features match, their product can be large regardless of sign, whereas summation yields large activations if either input is large. We generalize this idea by predicting affine transforms that are applied to the features of $t$ and $r$, followed by a sigmoid; see Fig. S5 (bottom). Two per-pixel, per-channel masks are thus predicted, $\mathbb{I}_{t}$, $\mathbb{I}_{r}$. Errors are corrected by a joint mask predictor that inspects both $\mathbb{I}_{t}$, $\mathbb{I}_{r}$ (see Sec. E.2). Masks $\mathbb{I}_{t}$ and $\mathbb{I}_{r}$ are resampled 2×2$\times$2 × and multiplied with $\phi_{M}$ to project its features into subspaces for $T$, $R$. This key step assumes that the identity $\mathbb{I}_{t}$, $\mathbb{I}_{r}$ of the component to which each feature belongs is low in spatial frequency. By resampling masks, not features, sharp features are preserved. Errors are corrected with finishing convolutions, which render $T,R$.
Training uses a cycle-consistency loss, losses similar to [35], and begins from scratch. See Sec. E.2 for details.
@@ 5 Results
We evaluate the simulation, base model, and upsampler; extensive results are added in the supplement. We make four contributions. First, we show that dereflection models that are trained solely on simulated reflections can generalize to real images without fine-tuning on real images, provided that the simulation uses RAW images in a photometrically accurate way. Second, training and testing on RAW images improves results significantly, and more than prior model variations. Third, a context image can disambiguate the reflection content if it is captured in another direction (e.g., the selfie). Fourth, upsampling low-res results, which is imperative but largely neglected in the literature, works better if one explicitly matches features in the low-res outputs $t,r$ to the low-res mixture, and masks them from the high-res mixture to recover the high-res transmission and reflection.
@@ 5.1 Reflection simulation
Figure 3: The importance of synthesizing training data (top row) from linear images (middle row), compared to prior work. (a) Photometrically accurate illuminant colors are simulated by mixing before white balancing; mixing 8-bit white balanced images is much different. (b) Mixing in scene-referred linear units produces reflections that are strong in the shadows, but transparent in the highlights. (prior work) Such effects are visibly incorrect in prior work, which blend $8$-bit tone mapped images [52, 11]. (bottom) Real and simulated examples are shuffled together. For each real image, a similar synthetic reflection was manually found in the dataset. Real images were not captured to match known examples; these qualitative matches exist because the dataset size exceeds $10^{6}$ (even numbered images are synthetic) .
Figure 4: Results at $2048$p; base outputs inset.
Figure 5: Upsampling GT images $256$p to $2048$p. V-DESIRR [35] adds artifacts.
@@
Source images were drawn from MIT5K [8], RAISE [10], and Laval Indoors [14], totaling $12$,$803$ RAWs and $2$,$233$ scene-referred Image-Based Lighting (IBL) panoramas. The $360^{\circ}$ IBLs are equivalent to about $12$,$367$ indoor RAW images because we simulate random cameras with an average FOV of $65^{\circ}$, Sec. B.2, B.4. Images are grouped into $10$,$547$ outdoor and $14$,$623$ indoor to create pairs $(i,j)$, Sec. D.2. The groups are split into train, validation, and test sets (80%, 15%, 5%) before simulation.
@@
The number of examples $(m,t,r)$ is amplified by randomization in the geometric simulation (Appendix B, D). We search $10^{8}$ examples for useful $m$. After culling, about $10^{7}$ mixtures remain, and we rendered $10\%$ at $256^{2}$ and $2048^{2}$ pixels to train the models. The $256^{2}$ pixel dataset has $1$,$241$,$091$ for training, $46$,$121$ for test, and $8$,$991$ for validation; the $2048^{2}$ dataset has $1$,$079$,$631$; $39$,$916$; and $7$,$448$.
@@
Fig. 5 shows results of mixing scene-referred images: (a) correct illuminant colors and (b) correct reflection visibility. We linearly blended $8$-bit tone-mapped images for comparison, and compare to prior works (see caption). Fig. S2 shows an overview of the dataset, and is discussed below.
In Tab. 3 we ablate each simulation component. We (gamma) compressed $t$, $r$ before compositing; separately exposed them (exposure); did not constrain their inclination or field-of-view (pose); removed spatially varying attenuation by making all camera rays normal to the glass (fresnel); separately white balanced (WB); removed depth-of-field blur and double reflection (blur); and removed all. We trained our base model on each dataset, and evaluated on the full-simulation test set. Each feature affects performance (all differences are significant), and omitting them all (All $\pm$ G) decreases performance dramatically compared to the average degradation to $t,r$ (control).
@@
Discussion. Prior works mix $8$-bit tone-mapped images, and the results are qualitatively unrealistic. Their simulated reflections overpower the highlights, and are not powerful enough in the shadows, which are boosted by tone mapping. In our accurate simulations, light from two scenes is mixed linearly and equally without tone-mapping. This accurate mixing allows our models to generalize better to new scenes, and yields SOTA performance without training on real images (Sec. 5.2). Furthermore, by synthesizing physically accurate reflections and searching for visible ones, we introduce natural priors on their appearance. Indoor light is weak, so reflections of the indoors are typically of regions near light sources or windows; see Fig. S2, examples $1$, $2$, $5$, $11$, $14$, $19$, $24$, $25$. Indoor lights create small reflections that often look yellow atop outdoor scenes, due to typical illuminant colors, whereas outdoor light can bounce off diffuse objects with enough strength to create colorful reflections of whole scenes that can be blue in white balance due to the outdoor illuminant color; see Fig. S2, examples $4$, $8$, $10$, $12$, $15$, $16$, $17$, $18$, $21$. At dusk, whole indoor scenes can reflect over cityscapes, etc. (examples $3$, $13$). Such priors are apparent in consumer photos (see Figs. 1, 8, 8, S6, S7, S8). Lastly, like prior works we pair indoor/outdoor photos, which permits pairings such as bathrooms and beaches. Such pairs can be removed if they prove unhelpful.
@@ 5.2 Base reflection removal
We trained three base models, one with and two without context $c$. To omit $c$, we removed the modulated merges (Fig. S4), which decreases model capacity. As a second option, we left the model unchanged, and trained/tested with random $c$. We used this second approach for ablation.
Method $\mathrm{SSIM}_{\mathrm{t}}$$\uparrow$ $\mathrm{SSIM}_{\mathrm{r}}$$\uparrow$
Control $85.8$ $49.0$
Full Sim. $\mathbf{95.7}$ $\mathbf{70}$ $\mathbf{88.2}$ $\mathbf{77}$
Ablate (G)amma $88.8$ $21$ $77.0$ $55$
Exposure $94.2$ $59$ $86.1$ $73$
Pose $95.0$ $65$ $85.4$ $71$
Fresnel $95.3$ $67$ $87.3$ $75$
WB $95.5$ $68$ $87.6$ $76$
Blur $95.5$ $68$ $87.9$ $76$
All -- G $90.2$ $31$ $71.3$ $44$
All $+$ G $89.2$ $24$ $58.5$ $19$
Table 1: Ablated datasets were created, $1.2$M examples each. Model ours+ctx was trained on these, and tested on the full-simulation test set. $\uparrow$ is w.r.t. control (see Tab. 3). SSIM values are shown as percentages.
Method $\mathrm{PSNR}_{\mathrm{t}}$ $\mathrm{SSIM}_{\mathrm{t}}$$\uparrow$ $\mathrm{PSNR}_{\mathrm{r}}$ $\mathrm{SSIM}_{\mathrm{r}}$$\uparrow$
Control $21.74$ $85.8$ $12.48$ $49.0$
RAW Train Ours+ctx $\mathbf{33.23}$ $\mathbf{95.7}$ +$\mathbf{70}$ $\mathbf{30.17}$ $\mathbf{88.2}$ +$\mathbf{77}$
Ours $32.15$ $95.2$ +$66$ $29.18$ $86.7$ +$74$
Zhu [60] $29.84$ $92.8$ +$49$
DSRNet [20] $28.98$ $92.6$ +$48$ $23.99$ $75.5$ +$52$
Zhang [59] $26.23$ $89.9$ +$28$ $22.78$ $61.5$ +$25$
CoRRN [47] $22.75$ $86.7$ +$6$ $18.31$ $60.4$ +$22$
$8$-bit Pub. Control $18.62$ $78.4$ $9.79$ $37.4$
DSRNet [20] $19.99$ $80.0$ +$8$ $16.98$ $49.3$ +$19$
Zhu [60] $19.84$ $79.7$ +$6$
Zhang [59] $18.65$ $75.9$-$11$ $17.37$ $51.0$ +$22$
CoRRN [47] $18.95$ $74.7$-$17$ $15.99$ $23.8$-$22$
Abl. Ours+rac $33.20$ $95.7$ +$70$ $30.20$ $88.3$ +$77$
Ours+rnd $32.42$ $95.1$ +$66$ $29.29$ $86.7$ +$74$
Table 2: Base models: (control) compares $m$ to $t$, $r$. $8$-bit models use published weights. $\uparrow$ is w.r.t. SSIM control. Ablations: (rac) GT $r$ is used as context; (rnd) random $c$. Ours+ctx beats Ours+rnd ($p<1.7e$-$11$).
@@
Control $19.50$ $86.3$ $13.13$ $64.5$
GT Ours $\mathbf{47.77}$ $\mathbf{98.8}$ $\mathbf{45.93}$ $\mathbf{98.2}$
Ours-NM $43.29$ $98.1$ $43.99$ $96.9$
VDSR+C [35] $42.24$ $97.9$ $38.32$ $93.8$
VDSR [35] $40.74$ $97.4$ $38.30$ $93.9$
Bicubic $31.98$ $85.1$ $41.58$ $96.0$
SUPIR [58] $28.09$ $64.6$ $28.29$ $56.8$
RESRGan [50] $23.72$ $65.6$ $23.11$ $53.6$
E2E Ours $\mathbf{30.62}$ $\mathbf{95.2}$ $\mathbf{28.53}$ $\mathbf{90.7}$
VDSR+C [35] $30.27$ $94.5$ $27.74$ $88.7$
@@
Table 3: Upsampling ground truth (GT), and using the base model for end-to-end results (E2E). Usampling is from $256$p to $2048$p using our method and V-DESIRR with and without cycle consistency (+C), which improves VDSR for $\mathrm{T}$ ($p$$<$$1e$-$12$).
Our system uses RAW images end-to-end, but public datasets do not provide RAW images: Real20, Real45, Nature, SIR2, SIR2+, CDR, [^2] The authors of CDR [28] have not released the RAW data. and RRW all use JPG/PNG formats [59, 11, 29, 46, 48, 28, 60]. We tabulate results using our simulation test sets, and show visual results using RAW photos that were captured in-the-wild. See also Sec. F.2.
In Tab. 3 we compare to Zhang et al. [59], DSRNet [20], Zhu et al. [60], and CoRRN [47] by retraining their models on our RAW dataset. [^3] We use $2.5$M parameters; DSRNet uses $125$M. Inference at $256\times 341$ takes $0.96\mathrm{s}$/$1.04\mathrm{s}$ on a $2021$ M1 MacBook Pro ($32$Gb) and iPhone $14$ Pro. Recall that our model uses the same losses and network head as Zhang et al. This simplifies comparison to prior work. Tab. 3 (RAW Train) shows that, when training with RAW, all methods improve images relative to the average degradation to $t,r$ (control). Our models however outperform prior works (ours+ctx, ours).
[…]
@@
In Tab. 3 (Abl.) we ablate our contextual model by training/testing with random $c$ (ours+rnd), which degrades performance compared to (ours+ctx)—this is statistically significant. Removing operations that use $c$ (ours) did not degrade performance compared to ours+rnd ($p<1.7e$-$11$), which suggests that ours+rnd does not learn dataset priors with its additional capacity, and conversely that ours+ctx leverages the content of $c$. Ablating further, using the reflection as the context ($c=r$) at test time only does not improve the contextual model results (ours+rac), which suggests that $c$ and $r$ need not match; the model is robust to their differences since it is trained with disjoint crops ($c$, $r$).
@@
For visual comparison, in Fig. 8 and Fig. S6 we captured [^5] We thank Florian Kainz for his help capturing these photos. ground truth reflections in common cases: looking outdoors, into a display case, and at artwork. We dereflected with Zhang, DSRNet (retrained, RT), and our models at $256\times 384$ (inset images) and upsampled to $2048\times 2731$ (next section). The empirical SSIM values (lowercase t, r) are commensurate with test performance (Tab. 3). In Fig. 8 our contextual model separates the reflection, but without context our model attributes the colors in the umbrella with a reflected object. Prior works perform quantitatively worse.
@@
In Figs. 1, 5, 8, S7, S8, and S9 we show results on photos in-the-wild from cameras that were not used to construct the training data. We also compare the $8$-bit models of Zhang and DSRNet. The bottom two rows of Figs. 8, S7, S8 show that these prior $8$-bit models perform qualitatively worse than when they are re-trained/evaluated on RAW (the top rows). They do not however recover $r$ well, which is needed for aesthetics and error correction (Sec. 5.4, Sec. F.4).
@@
Discussion. Our models recover $t,r$ in diverse real-world cases including museums, nature, shopping, a mid-day city, artwork, etc. (Figs. 1, 5, 8, 8, S6, S7, S8, S9). In Fig. 1, using the context photo yields more correct and uniform color on the Egyptian tablet because there is less ambiguity about the color of the reflection scene (compare to inset w/o context). Failures occur when $t$ or $r$ is bright, and pushes the other into the noise floor, saturating it to black—the problem becomes hole filling. When a single color channel saturates, the content can sometimes be recovered. But, systems must address hole filling because users typically cannot control the strength of reflections.
@@
Errors can occur when textured regions of $t$ and $r$ overlap, as in Fig. 1 where a stone wall overlaps the subject’s dress. Color differences help: in Fig. 8 the reflected painting is separated from the tree. Without such differences, models must repair or hallucinate content in the corrupted $t$. Saturated reflections pose a similar challenge. See Appendix F for more discussion of errors and additional results.